# Are Ensembles Getting Better All the Time?

### v2

@@ 4.3 Monotonicity of tail probabilities
Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{D}$ be i.i.d. random vectors distributed as $X$, finite mean $\mu$, moment-generating function $\phi:t\mapsto\operatorname{\mathbb{E}}\left[\mathrm{exp}\left(\langle t,X_{1}\rangle\right)\right]$, and cumulant-generating function $\varphi=\log\phi$. Assume that
@@
    $X$ has density $\rho$ with respect to the Lebesgue measure on $\mathbb{R}^{D}$ and that $\rho$ does not cancel;
@@
    there exists $\alpha>0$ such that $\phi$ is differentiable and bounded on $U_{\alpha}$, the open ball of radius $\alpha$
@@
$\min_{i\leq D}\frac{\varepsilon_{i}}{\left\lVert\varepsilon\right\rVert}<\sqrt{\frac{\lambda_{\min}}{\lambda_{\max}}}\,,$ (49)
@@ 4.4 Ensembles and the classification error
We are finally in position to deal with the classification error. We remind the notations introduced in Section 4.2.3. We deal with $n_{\text{Cl}}$ classes, and the true label, seen as a 1-hot encoding, is $y\in\{0,1\}^{n_{\text{Cl}}}$. A prediction is a $n_{\text{Cl}}$-dimensional vector $\hat{y}=(\hat{y}^{(c)})_{c\leq n_{\text{Cl}}}\in\mathbb{R}^{n_{\text{Cl}}}$ of scores. Without loss of generality, we assume that the true class is the first one: $y=(1,0,\ldots,0)$. Our goal is to use the results of the previous section together with the following representation representation, obtained in Lemma 6:
@@
In the binary case, when $n_{\text{Cl}}=2$, $X_{k}={\hat{y}_{k}}^{(1)}-{\hat{y}_{k}}^{(2)}$ just corresponds to the margin, that is, the difference between the score of the correct label and the score of the incorrect one. In the multiclass case, $X_{k}$ is a $(n_{\text{Cl}}-1)$-dimensional vector of margins obtained for each possible “1 versus $c$” classification problems, for $c\in\{2,\ldots,n_{\text{Cl}}\}$. Following the seminal work of Bartlett et al. (1998), margins have been used a lot to study the theory of ensembles (see, e.g., Biggs et al., 2022, and references therein).
@@
[…]
Assumption 1 implies in particular that $\varepsilon=\mathbb{E}[X_{1}]>0$ which is equivalent to the fact that the expected margin ${\overline{y}_{\infty}}^{(1)}-\max_{c\neq 1}{\overline{y}_{\infty}}^{(c)}$ is positive, and to the fact that the asymptotic prediction is correct, that is, $L(\overline{y}_{\infty})=0$. This assumption of positive margin is often used to study ensembles (see, e.g., Breiman, 2001).
@@
[…]
@@
[…]
@@
Proof  We begin by assuming that the prediction is correct (Assumption 1). Then, Theorem 7 (when $n_{\text{Cl}}=2$) or Theorem 8 (when $n_{\text{Cl}}>2$) ensures that $\operatorname{\mathbb{P}}\left(-\overline{X}_{n}\geq-\mathbb{E}[X_{1}]+\varepsilon\right)=\operatorname{\mathbb{P}}\left(-\overline{X}_{n}\geq 0\right)$ decreases for $K$ large enough. Using Lemma 6, this implies that $L(\overline{y}_{\infty})$ eventually decreases.
We now deal with the case of an incorrect prediction (Assumption 2). Again, Theorem 7 or Theorem 8 ensures that $\operatorname{\mathbb{P}}\left(\overline{X}_{n}>\mathbb{E}[X_{1}]+\varepsilon\right)=\operatorname{\mathbb{P}}\left(\overline{X}_{n}>0\right)$ eventually decreases. In turn, $L(\overline{y}_{\infty})$ eventually increases.
@@ 5.1 Classification of skin lesions with convolutional networks
We start by detailing the motivating example of Section 1.1. We want to predict whether a dermatoscopic image corresponds to a benign keratosis-like lesion or a melanoma. We use the DermaMNIST (Yang et al., 2023) data set, based on the HAM10000 collection (Tschandl et al., 2018), and retain only the classes “benign keratosis” and “melanoma,” because they are balanced within DermaMNIST (the whole dataset is very unbalanced) and notoriously difficult to distinguish (Grant-Kels et al., 1999). The training/validation/test split is the same as the one from Yang et al. (2023), and consists of $1,548$ color images of resolution $28\times 28$ for training, $221$ similar images for validation, and $443$ test images.
@@ 5.2 Wisdom of crowds for movie rating predictions
Figure 5: Evolution of the error at predicting the ratings of three movies as the size of the crowd grows (mean and standard deviation over $10,000$ repetitions). (Left) The squared error is decreasing for all three movies. Since this is a convex loss, this is in line with Theorem 2. (Middle and Right) The Welsh and Geman-McClure losses decreases for Ghost in the Shell, but increases for the other two movies. This is in line with Theorem 5, that implies that, for these two losses, the loss should be eventually decreasing when $|y-\bar{y}_{\infty}|<10$, and increasing when $|y-\bar{y}_{\infty}|>10$.
@@ 6 Conclusion
Our initial question was “Is it always true that an ensemble of $K+1$ models performs better than an ensemble of $K$ models?” When the loss is convex and the ordering of the ensemble does not matter, Theorems 2 and 4 provide a clear and positive answer to this question. When the loss is non-convex, however, a more nuanced picture must be drawn: ensembles will get better in some setting, and worse in others.
@@
This work was supported by the French government, through the 3IA Côte d’Azur Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-19-P3IA-0002. DG also acknowledges the support of ANR through project NIM-ML (ANR-21-CE23-0005-01) and of EU Horizon 2020 project AI4Media (contract no. 951911).
@@ Appendix A Proof of strict improvements under strong convexity
$\small L\left(\frac{1}{K}\sum_{k=1}^{K}x_{k}\right)\leq\frac{1}{K}\sum_{k=1}^{K}L\left(x_{k}\right)-\mu\frac{1}{K}\sum_{k=1}^{K}\left|\left|x_{k}-\frac{1}{K}\sum_{j=1}^{K}x_{j}\right|\right|^{2}.$ (53)
@@
$\mathbb{E}\left[L\left(\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}\right)\right]\leq\mathbb{E}\left[L\left(\frac{1}{K-1}\sum_{k=1}^{K-1}\hat{y}_{k}\right)\right]-\mu\mathbb{E}\left[\left|\left|\frac{1}{K-1}\sum_{k=1}^{K-1}\hat{y}_{k}-\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}\right|\right|^{2}\right].$ (54)
@@
$\mathbb{E}\left[\left|\left|\frac{1}{K-1}\sum_{k=1}^{K-1}\hat{y}_{k}-\frac{1}{K}\sum_{k=1}^{K}\hat{y}_{k}\right|\right|^{2}\right]$ $=\frac{1}{K^{2}(K-1)^{2}}\mathbb{E}\left[\left|\left|K\sum_{k=1}^{K-1}\hat{y}_{k}-(K-1)\left(\hat{y}_{K}+\sum_{k=1}^{K-1}\hat{y}_{k}\right)\right|\right|^{2}\right]$
$=\frac{1}{K^{2}(K-1)^{2}}\mathbb{E}\left[\left|\left|\sum_{k=1}^{K-1}\hat{y}_{k}-(K-1)\hat{y}_{K}\right|\right|^{2}\right]$
@@
Appendix B More details on the spherical score
@@
when the true label is $y=1$. When the true label is 00, as in Section 4.1, it is defined as
@@
The second derivative of $L_{0}$ is $\hat{y}\mapsto(-4\hat{y}^{2}+\hat{y}+1)/(2(\hat{y}-1)x+1)^{5/2}$, whose unique root in $[0,1]$ is $1/8+\sqrt{17}/8\approx 0.64$. Therefore, as seen in Figure 3, $L_{0}$ will be strictly convex on $[0,1/8+\sqrt{17}/8)$ and strictly concave on $(1/8+\sqrt{17}/8,1]$.
@@ Appendix C Proof of the theorem for smooth losses (Theorem 5)
We only treat the case $H(\hat{y}_{\infty})\succ 0$, since the case $H(\hat{y}_{\infty})\prec 0$ follows by replacing $L$ by $-L$. We will base our proof on the following lemma, which is obtained by Taylor-expanding the loss up to the fourth order. In this context, such a technique is often called the delta method for moments (see, e.g., Small, 2010; Bickel and Doksum, 2015). The proof of the lemma is slightly tedious, and is provided after the main theorem.
@@
Let $X_{1},\ldots,X_{n}$ be a sequence of i.i.d. random variables whose first $5$ moments are finite, with mean $\mu$ and covariance matrix $\Sigma\succeq 0$. Let $L$ be a function with continuous and bounded partial derivatives of order up to 5, with Hessian matrix $H$. Then, there exists a constant $\alpha\in\mathbb{R}$, that depends on the first $3$ moments of $X$ and the partial derivatives of the loss of order up to $4$, such that, when $K$ goes to infinity
@@
Applying this result successively to $\bar{y}_{K}$ and $\bar{y}_{K+1}$ and taking the difference between the two terms, we obtain

### v3

@@ 4.3 Monotonicity of tail probabilities
Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{D}$ be i.i.d. random vectors distributed as $X$ with finite mean $\mu$, moment-generating function $\phi:t\mapsto\operatorname{\mathbb{E}}\left[\mathrm{exp}\left(\langle t,X_{1}\rangle\right)\right]$, and cumulant-generating function $\varphi=\log\phi$. Assume that
@@
    $X$ has strictly positive density $\rho$ with respect to the Lebesgue measure on $\mathbb{R}^{D}$;
@@
    there exists $\alpha>0$ such that $\phi$ is differentiable and bounded on $U_{\alpha}$, the open ball of radius $\alpha$;
@@
$\min_{1\leq i\leq D}\frac{\varepsilon_{i}}{\left\lVert\varepsilon\right\rVert}<\sqrt{\frac{\lambda_{\min}}{\lambda_{\max}}}\,$, (24)
@@
Remark 9 (Monotonicity of tails under Gaussian assumption).
Let us first notice that the result of Theorem 8 is trivially true under the assumption that $X_{1},\ldots,X_{n}$ are i.i.d. $\mathcal{N}(0,\Sigma)$. Indeed, in this simplified setting,
$\operatorname{\mathbb{P}}\left(\overline{X}_{n}\geq\mu+\varepsilon\right)=\operatorname{\mathbb{P}}\left(\mathcal{N}(0,\Sigma)\geq\sqrt{n}\varepsilon\right)\,$,
which is decreasing since we integrate with respect to a positive density.
Nevertheless, let us check the assumptions of Theorem 8 in the multivariate Gaussian setting in order to show that they are not too stringent. First, the density of the multivariate Gaussian does not cancel, ensuring that Assumption (1) is met. Second, it is known that
$\forall t\in\mathbb{R}^{D},\qquad\phi_{X}(t)=\exp\left(\mu^{\top}t+\frac{1}{2}t^{\top}\Sigma t\right)\,$,
which is bounded and differentiable on any open ball of finite radius, yielding Assumption (2). Third, we observe that $\nabla^{2}\phi=\Sigma$, a positive-definite matrix. Henceforth Assumption (3) is satisfied.
Then remains the question of the existence of $\varepsilon\in\mathbb{R}^{D}$ such that $\varepsilon>0$ and Eq. (24) holds—after all, this condition could be impossible to meet, even in such a simple scenario. We see that it is not the case: taking $\lambda_{\min}=\lambda_{\max}=1$ gives rise to the condition $\min_{i}\varepsilon_{i}^{2}<\left\lVert\varepsilon\right\rVert^{2}$, which is true for any $\varepsilon\in\mathbb{R}^{D}$ provided that $D\geq 2$. To see this, one can reason by contradiction and suppose that $\min_{i}\varepsilon_{i}\geq\left\lVert\varepsilon\right\rVert$, meaning that for all $i\in[D]$, $\varepsilon_{i}^{2}\geq\left\lVert\varepsilon\right\rVert^{2}$. Summing these inequalities returns $\left\lVert\varepsilon\right\rVert^{2}\geq D\left\lVert\varepsilon\right\rVert^{2}$, a contradiction. A similar proof shows that any ratio $\lambda_{\min}/\lambda_{\max}\geq 1/D$ does not actually put a constraint on $\varepsilon$.
When the $\lambda_{\min}/\lambda_{\max}$ is smaller, the condition over $\varepsilon$ becomes more stringent, but the set of $\varepsilon$ remains non-empty. Indeed Eq. (24) one can always take one of the coordinates of $\varepsilon$ to be arbitrarily small with respect to the others, automatically satisfying the condition. For instance, one can consider $\varepsilon$ such that $\varepsilon_{1}=\sqrt{\lambda_{\min}}$ and all other coordinates equal to $\sqrt{\lambda_{\max}}$. Moreover, we note that Eq. (24) is a cone condition, meaning that all (positive) multiples of an $\varepsilon$ satisfying the condition satisfies it itself. In definitive, Theorem 8 holds true for any multivariate Gaussian distribution, with a set of $\varepsilon$ being more and more constrained, but never empty.
@@ 4.4 Ensembles and the classification error
We are finally in position to deal with the classification error. We remind the notations introduced in Section 4.2.3. We deal with $n_{\text{Cl}}$ classes, and the true label, seen as a 1-hot encoding, is $y\in\{0,1\}^{n_{\text{Cl}}}$. A prediction is a $n_{\text{Cl}}$-dimensional vector $\hat{y}=(\hat{y}^{(c)})_{c\leq n_{\text{Cl}}}\in\mathbb{R}^{n_{\text{Cl}}}$ of scores. Without loss of generality, we assume that the true class is the first one: $y=(1,0,\ldots,0)$. Our goal is to use the results of the previous section together with the following representation, obtained in Lemma 6:
@@
In the binary case, when $n_{\text{Cl}}=2$, $X_{k}={\hat{y}_{k}}^{(1)}-{\hat{y}_{k}}^{(2)}$ just corresponds to the margin, that is, the difference between the score of the correct label and the score of the incorrect one. In the multiclass case, $X_{k}$ is a $(n_{\text{Cl}}-1)$-dimensional vector of margins obtained for each possible “1 versus cc” classification problems, for $c\in\{2,\ldots,n_{\text{Cl}}\}$. Following the seminal work of Bartlett et al. (1998), margins have been used a lot to study the theory of ensembles (see, e.g., Biggs et al., 2022, and references therein).
@@
Let $\varepsilon=\mathbb{E}[X_{1}]={\overline{y}_{\infty}}^{(1)}1_{n_{\text{Cl}}-1}-{\overline{y}_{\infty}}^{(-1)}$, where we set $u^{(-1)}=(u_{2},\ldots,u_{p})^{\top}$ for any vector of $\mathbb{R}^{p}$. There are either $n_{\text{Cl}}=2$ classes and the regularity conditions of Theorem 7 are verified, or there are $n_{\text{Cl}}>2$ classes and the regularity conditions of Theorem 8 are verified, for $X_{1},\ldots,-X_{K}$ and $\varepsilon$.
Assumption 1 implies in particular that $\varepsilon=\mathbb{E}[X_{1}]>0$ which is equivalent to the fact that the expected margin ${\overline{y}_{\infty}}^{(1)}-\max_{c\neq 1}{\overline{y}_{\infty}}^{(c)}$ is positive. This assumption of positive margin is often used to study ensembles (see, e.g., Breiman, 2001). It implies, in particular, that the asymptotic prediction is correct, that is, $L(\overline{y}_{\infty})=0$.
@@
Let $\varepsilon=-\mathbb{E}[X_{1}]={\overline{y}_{\infty}}^{(-1)}-{\overline{y}_{\infty}}^{(1)}1_{n_{\text{Cl}}-1}$. There are either $n_{\text{Cl}}=2$ classes and the regularity conditions of Theorem 7 are verified, or there are $n_{\text{Cl}}>2$ classes and the regularity conditions of Theorem 8 are verified, for $X_{1},\ldots,X_{K}$ and $\varepsilon$.
@@
Let $\hat{y}_{1},\ldots,\hat{y}_{K}\in\mathbb{R}^{n_{\text{Cl}}}$ be i.i.d. random variables. Then
@@
Proof We begin by assuming that the prediction is correct (Assumption 1). Then, Theorem 7 (when $n_{\text{Cl}}=2$) or Theorem 8 (when $n_{\text{Cl}}>2$) ensures that $\operatorname{\mathbb{P}}\left(-\overline{X}_{n}\geq-\mathbb{E}[X_{1}]+\varepsilon\right)=\operatorname{\mathbb{P}}\left(-\overline{X}_{n}\geq 0\right)$ decreases for $K$ large enough. Using Lemma 6, this implies that $L(\overline{y}_{K})$ eventually decreases.
We now deal with the case of an incorrect prediction (Assumption 2). Again, Theorem 7 or Theorem 8 ensures that $\operatorname{\mathbb{P}}\left(\overline{X}_{n}>\mathbb{E}[X_{1}]+\varepsilon\right)=\operatorname{\mathbb{P}}\left(\overline{X}_{n}>0\right)$ eventually decreases. In turn, $L(\overline{y}_{K})$ eventually increases.
@@ 5.1 Classification of skin lesions with convolutional networks
We start by detailing the motivating example of Section 1.1. We want to predict whether a dermatoscopic image corresponds to a benign keratosis-like lesion or a melanoma. We use the DermaMNIST (Yang et al., 2023) data set, based on the HAM10000 collection (Tschandl et al., 2018), and retain only the classes “benign keratosis” and “melanoma”, because they are roughly balanced within DermaMNIST (the whole dataset is very unbalanced) and notoriously difficult to distinguish (Grant-Kels et al., 1999). The training/validation/test split is the same as the one from Yang et al. (2023), and consists of $1,548$ color images of resolution $28\times 28$ for training, $221$ similar images for validation, and $443$ test images.
@@
To also illustrate our results in the multiclass setting, we consider a version of the same dataset with an additional class, corresponding to basal-cell carcinoma, which is the most common type of skin cancer. We present our results in Figure 5. Here, while the assumptions of Theorem 10 are difficult to check in practice, the conclusion of the theorem seems to hold: the accuracy increases for correct predictions and decreases for completely incorrect predictions. For predictions that are neither, the accuracy is empirically decreasing. We conjecture that some extension of Theorem 10 should explain this phenomenon.
Figure 5: Evolution of the accuracy of a dropout ensemble on the 3-class dermatology data set as the number of models grows (mean and standard deviation over $500$ repetitions). (Left) The accuracy averaged over the whole test set has a non-monotonic behaviour. (Centre-left) The accuracy averaged over images for which the prediction is asymptotically correct is improving. (Centre-Right) The accuracy averaged over images for which the prediction is asymptotically incorrect (yet the correct label is not the one with the smallest predicted probability) is degrading. (Right) The accuracy averaged over images for which the prediction is asymptotically completely incorrect (i.e. the correct label has the smallest predicted probability) is degrading. Theorem 10 provides a theoretical explanation of the behaviour of centre-left and right panels.
@@ 5.2 Wisdom of crowds for movie rating predictions
Figure 6: Evolution of the error at predicting the ratings of three movies as the size of the crowd grows (mean and standard deviation over $10,000$ repetitions). (Left) The squared error is decreasing for all three movies. Since this is a convex loss, this is in line with Theorem 2. (Middle and Right) The Welsh and Geman-McClure losses decreases for Ghost in the Shell, but increases for the other two movies. This is in line with Theorem 5, which implies that, for these two losses, the loss should be eventually decreasing when $|y-\bar{y}_{\infty}|<10$, and increasing when $|y-\bar{y}_{\infty}|>10$.
@@ 6 Conclusion
Our initial question was “Is it always true that an ensemble of $K+1$ models performs better than an ensemble of $K$ models?” When the loss is convex and the ordering of the ensemble does not matter, Theorems 2 and 4 provide a clear and positive answer to this question. When the loss is non-convex, however, a more nuanced picture must be drawn: ensembles will get better in some settings, and worse in others. Namely, ensembles will keep getting better (respectively worse) when their asymptotic prediction falls in a locally convex (respectively concave) spot of the loss.
6.1 Practical guidelines
Our work leads us to strengthen the practical guidelines provided by Probst and Boulesteix (2018, Section 5.2): in general, ensembles should be as large as computationally feasible. Indeed, we showed that this is always true for convex losses. Furthermore, even for nonconvex losses, it is likely that, when averaged over the whole dataset, the loss will be decreasing, since it is reasonable to expect to see more “good” points than “bad” points.
An additional empirical guideline is related to evaluation of ensemble techniques. When a new technique is proposed, we recommend to report both convex and nonconvex losses when it is possible. Indeed, since any ensemble will get better for a convex loss, showing such improvements will hardly be useful to assess the quality of an ensemble technique. Adding other nonconvex metrics can paint a more detailed picture. Concretely, when evaluating ensembles of classifiers, we recommend to always show, at the very least, the cross-entropy (and/or the Brier score) and the classification error. This guideline is already followed quite often (e.g. by Lakshminarayanan et al., 2017, Figure 2)
6.2 Future work
@@
All our results were at the individual level, and can be generalised to losses averaged over datasets using the fact that averages of decreasing sequences are decreasing. However, this reasoning will not work for metrics that are not simple averages. For instance, it would be interesting to study the area under the receiver operating characteristic curve or the median squared and errors. Probst and Boulesteix (2018) noticed non-monotonic behaviours for these three losses.
Finally, it would be interesting to assess whether or not ensembling can improve model calibration. This question was investigated by researchers working on scoring rules. In particular, it was shown that ensembles of calibrated models can be uncalibrated (Hora, 2004; Ranjan and Gneiting, 2010; Gneiting and Ranjan, 2013). It would be interesting to study under which conditions calibration may be monotonically improving.
@@
This work was supported by the French government, through the 3IA Côte d’Azur Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-19-P3IA-0002. DG also acknowledges the support of ANR through project NIM-ML (ANR-21-CE23-0005-01) and of EU Horizon 2020 project AI4Media (contract no. 951911). Parts of this work were done as DG was employed at Université Côte d’Azur. The authors thank Raphaël Razafindralambo for spotting an error in an earlier version of this paper.
@@ Appendix A Proof of strict improvements under strong convexity
[…]
@@
[…]
@@
[…]
@@
Appendix B More details on nonconvex losses
B.1 More details on the spherical score
@@
when the true label is $y=1$. When the true label is 0, as in Section 4.1, it is defined as
@@
The second derivative of $L_{0}$ is $\hat{y}\mapsto(-4\hat{y}^{2}+\hat{y}+1)/(2(\hat{y}-1)x+1)^{5/2}$, whose unique root in $[0,1]$ is $1/8+\sqrt{17}/8\approx 0.64$. Therefore, as seen in Figure 3, $L_{0}$ will be strictly convex on $0,1/8+\sqrt{17}/8)$ and strictly concave on $(1/8+\sqrt{17}/8,1]$.
B.2 More details on the tangent loss
The tangent loss (Masnadi-Shirazi et al., 2010) is a loss function for binary classification. There are two labels, $y=1$ and y=1y=-1, and the output of our model is a score $\hat{y}\in\mathbb{R}$ that indicates its confidence in predicting that the label is $1$. When the true label is $y\in\{-1,1\}$, it is defined as
$L_{y}(\hat{y})=(2\textup{arctan}(y\,\hat{y})-1)^{2}.$ (35)
This function is convex around its minimum of one, but nonconvex when preditions are overconfident. Indeed, when $y=1$, the fonction is locally convex on $y_{-},y_{+}]$ with $y_{-}\approx-0.51$ and $y_{+}\approx 1.26$, and locally concave elsewhere.
@@ Appendix C Proof of the theorem for smooth losses (Theorem 5)
We only treat the case $H(\hat{y}_{\infty})\succ 0$, since the case $H(\hat{y}_{\infty})\prec 0$ follows by replacing $L$ by L-L. We will base our proof on the following lemma, which generalises a result of Small (2010, Section 4.3), and is obtained by Taylor-expanding the loss up to the fourth order. In this context, such a technique is often called the delta method for moments (see, e.g., Small, 2010; Bickel and Doksum, 2015). The proof of the lemma is slightly tedious, and is provided after the main theorem.
@@
Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{d}$ be a sequence of nondegenerate i.i.d. random variables, whose first $5$ moments are finite, with mean $\mu$ and covariance matrix $\Sigma\succeq 0$. Let $L$ be a function with continuous and bounded partial derivatives of order up to 5, with Hessian matrix $H$. Then, there exists a constant $\alpha\in\mathbb{R}$, that depends on the first $3$ moments of $X$ and the partial derivatives of the loss of order up to $4$, such that, when $K$ goes to infinity
@@
Proof of Theorem 5. Applying Lemma 12 successively to $\bar{y}_{K}$ and $\bar{y}_{K+1}$ and taking the difference between the two terms, we obtain